How many millliters of 0.250 M NaOH are required to neutralize the following solutions? a. 60.0 mL of 0.0750 MHC b. 35.0 mL of 0.226 MHNO C.
Answer #1

a)
Balanced chemical equation is:
HCl + NaOH ---> NaCl + H2O


Here:
M(HCl)=0.075 M
M(NaOH)=0.25 M
V(HCl)=60.0 mL

According to balanced reaction:
1*number of mol of HCl =1*number of mol of NaOH
1*M(HCl)*V(HCl) =1*M(NaOH)*V(NaOH)
1*0.075 M *60.0 mL = 1*0.25M *V(NaOH)
V(NaOH) = 18 mL

Answer: 18.0 mL

b)
Balanced chemical equation is:
HNO3 + NaOH ---> NaNO3 + H2O


Here:
M(HNO3)=0.226 M
M(NaOH)=0.25 M
V(HNO3)=35.0 mL

According to balanced reaction:
1*number of mol of HNO3 =1*number of mol of NaOH
1*M(HNO3)*V(HNO3) =1*M(NaOH)*V(NaOH)
1*0.226 M *35.0 mL = 1*0.25M *V(NaOH)
V(NaOH) = 31.64 mL

Answer: 31.6 mL

C)
Balanced chemical equation is:
H2SO4 + 2 NaOH ---> NaCl + 2 H2O


Here:
M(H2SO4)=0.19 M
M(NaOH)=0.25 M
V(H2SO4)=75.0 mL

According to balanced reaction:
2*number of mol of H2SO4 =1*number of mol of NaOH
2*M(H2SO4)*V(H2SO4) =1*M(NaOH)*V(NaOH)
2*0.19 M *75.0 mL = 1*0.25M *V(NaOH)
V(NaOH) = 114 mL

Answer: 114 mL